Status Report

ISS On-Orbit Status 12 Mar 2002

By SpaceRef Editor
March 12, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Crew performance continues to show unflagging excellence. This morning, Flight Control gratefully commended yesterday’s work by Dan Bursch to return the RED exerciser to full operations ahead of schedule.

Preparations for Robotics Day #4 (3/14) are underway, and troubleshooting continues on the ground regarding the SSRMS anomalies experienced last week. For now, problems can be bypassed by workarounds, but further decisions on returning the SSRMS to full operational capability on both electronics “strings” are pending. [As reported to the crew, there is a failure in the Wrist Roll Joint (when on the Prime String) which prevents release of the joint brakes. The Redundant String works OK. A “bug” has also been identified in the RWS OCS (robotics workstation Operations & Control Software), which locks up when the brake problem is annunciated. A s/w patch to correct this is in discussion. (The OCS was downloaded from the C&C MDM into the memory of the joint’s Êcontrol electronics unit, CEU, along with Video Graphics s/w, VGS, after the Workstation Host s/w, WHS, had been downloaded into the CEU). Two other s/w problems were encountered, viz., an address conflict between the Tip LEE (latching end effector) camera and Tip Elbow camera, and a switch-over failure from the Lab RWS to the CUPola RWS.] Ê

For the inspection/checkout of the Lab cradle assembly (LCA) by the ground via tip LEE camera, planned for tomorr|-#fternoon, the SSRMS wrist will be reconfigured to support the activity. (LCA is required for the S0 truss transfer during 8A). Also, additional brake testing will be performed on the Redundant string in order to gather baseline data, and the DOUG (dynamic operation ubiquitous graphics) setup will be reviewed by the crew for Robotics Day #4 on Thursday (3/14). On that day the SSRMS is required to complete some exterior surveys of ISS (radiator array panels and base plates for any surface discolorations or deformations [bubbles, flakes]), with only the Lab RWS and Redundant String used throughout the day. Time allowing, a fast LEE checkout is also desired once the survey is completed. The camera recordings will be downlinked later and the tape returned on 8A. On Friday (3/15), SSRMS cameras are required to provide viewing of the planned purge of the SM propellant lines, which were used to transfer props between FGB and Progress 6P. For this task, it is planned to use the CUPola RWS and Prime SSRMS string.

The R&R (remove & replace) activity at the Node starboard hatch was deferred to a later time to provide the ground more time to prepare an acceptable leak check technique required to verify hatch window integrity. The plan is to replace the current panel in the Node hatch to the Airlock (A/L) with a port window to allow crewmembers to look into the A/L when the hatch is closed. Today’s specific concern was that with the current procedure and leak criteria a leak in the window could be masked by known VAJ (vacuum access jumper) leakage.

The crew was uplinked instructions for assembling two ISS emergency kits for quick access in a time-critical contingency situation. These kits are now located in an easily accessible location in the Node (i.e., central place), and the crew would take them with them before closing hatches in case of a module evacuation. Two types of kits were prepared by Onufrienko, Walz and Bursch today: a “U.S. Post-Fire Kit” (two CTBs [crew transfer bags]), and an “ISS Leak Kit” (one CTB). The list of equipment contained in the kits was developed by ground specialists.

CDR Onufrienko replaced the IDZ-2 SPO smoke detector in the DC-1 docking compartment with another smoke detector (IDZ-2 SPOPT) from the FGB, in an effort to determine the optimal sensitivity threshold for the DS-1 (FGB smoke detectors have lower threshold).

The crew has gathered samples from different locations at the SM inner hull which have accumulated some mildew-like contaminant. A total of 10 samples were taken, six for Moscow and Êfour for Houston. Crew also reported that the substance appears to have formed in areas that had blocked airflow from stowage racks, and that there was no contamination in areas where the airflow was not blocked by stowage. The samples will be returned on both the Soyuz and 8A Shuttle for analysis. Subsequently, the crew also cleaned the areas that they could reach. Procedures are being developed for temporary removal of the TVIS treadmill (in its current configuration), to allow additional areas to be reached for cleaning.

CDR Onufrienko and FE-1 Walz continued loading operations on Progress M1-256, using the IMS (inventory management system) as reference and record “clipboard”. MCC-M has uplinked a long list of items to be transferred to the cargo ship as well as detailed instructions on how and where to stow discarded equipment and waste, in order to ensure approved CG (center-of-gravity) and mass characteristics (moments of inertia around three axes) necessary for the post-separation maneuvers of the automated vehicle. The waste cargo includes food waste bags, old medical kits, used filters and moisture collectors, discarded underwear and shoes from Expedition 3, pieces of hardware and Orlan gloves.

FE-2 Dan Bursch completed a series of CheCS (crew health care systems) hardware photographs for archival purposes. Images were taken of the CSA-CP (compound specific analyzer – combustion products), currently deployed in the SM, the ISS eyewash station near the SM SVO-ZV water system, the CheCS rack in the Lab, and the TEPC (tissue equivalent proportional counter) in the Lab temporary sleep station (TeSS), where Dan had installed it earlier this morning.

Photos taken on 3/8 of the ADVASC (Advanced Astroculture) growth chamber reveal that more plants are alive and developing than expected, although they are noticeably behind the schedule. Initial analysis indicates that a total of 14 plants are alive, and an additional 16 plants have germinated but did not survive. Investigators remain hopeful that other seeds will germinate before the end of the mission. Currently, the plan is to proceed with the plant tissue sample this Friday (3/15).

At 11:55 am EST, in sync with commands sent up by the ARIS-ICE (active rack isolation system/isolation characterization experiment) team, Carl Walz completed a critical performance test of the ARIS “snubber cups” containment. The test was to determine whether the ARIS rack, when moved up, hits the Lab standoff structure before the snubber pins hit the tops of the snubber cups.

Yuri Onufrienko returned to the previously conducted checkout/monitoring of the IK-0501 gas analyzer (GA) in the SM, taking control readings with the IGZ and BKGA gas meters. He also worked on recovering Russian Wiener laptop #1 functionality by copying the system file of laptop #2 to #1 via floppy disk.

All crewmembers performed their daily physical exercise on TVIS and the newly restored RED.

CEO (crew earth observation) target areas today were European Smog (this pass ran the length of the Italian Peninsula on the western side. Crew was to try for oblique and limb views of the favorable aerosol situation over eastern Italy and the Adriatic Sea to the left of track), W. Mediterranean Dust and Smog (with a strong cold front approaching the Straits of Gibraltar, crew was to look eastward, to the left of track for dust blowing northward out of Algeria and to use oblique or limb views only), Tropical Cyclone “Hary” (Dynamic Event Site: Tropical Cyclone Hary grazed the NE coast of Madagascar over the weekend with 140-kt winds before it weakened and headed south. This is still a major tropical cyclone with 115kt winds moving SSE over open water. Of interest this pass: near-nadir views of the eye and possible oblique views of the extent of the storm’s circulation), Congo-Zimbabwe Biomass Burning (the dry season is now underway for Zimbabwe and conditions are becoming favorable for burning. Crew was to look to the left of track for early indications of fires), Gulf of St. Lawrence (light is getting low this pass, but concentrate on ice coverage of the bay and coastal waters of Nova Scotia), Eastern United States (a slow-moving high-pressure area over the eastern US should have offered good conditions for viewing aerosols there. Of interest: looking to the right of track from the Lake Erie to the Delmarva Peninsula for layers and palls of smog that best detected in oblique views), and Lakes of the Eastern Sierra Watershed (snow pack in the Sierra Nevadas is nearing its seasonal peak for this year. Of interest this pass: documenting the extent of snow cover on these mountains just to the left of track).

Progress M1-256 (6P) thrusters are scheduled to perform another two-burn reboost tonight. The first burn (#5), at perigee, will start at 7:04 pm (5 min, delta-V about 2 m/s), the second (#6), at apogee, at 7:53 pm (10 min., 4 m/s). Attitude control has already been handed over to the Russian MCS (motion control system) at 3:00 pm EST, to be returned to CMG momentum management tonight at 8:30 pm.

Shuttle Columbia/STS-109, after an exceedingly successful mission to service and upgrade the Hubble Space Telescope, returned this morning to KSC at 4:32 am EST, after completing 163 orbits.

U.S. and Russian Segment Status (as of 1:48 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (32 Amps mode). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 760, temperature (deg C) — 26.5, ppO2 (mmHg) — 157.5, ppCO2 (mmHg) — 2.6.
  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) — 21.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 20.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 752.16, temperature (deg C) — 23.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 754.51, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 754.31, temperature (deg C) — n/a; shell heater temp (deg C) — 22.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 23.0.
  • PMA-2: Shell heater temp (deg C) — 16.7.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed (“parked”) position at 125 degrees.
  • SM batteries: battery #8 is cycling; all other batteries (7) in “partial charge” mode.
  • FGB battery #6 is offline; all other batteries (5) are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is back-up.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at PMA-3 clearance position (out of Soyuz emergency escape trajectory zone), with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 6:52 am EST):

  • Mean altitude — 385.2 km
  • Apogee — 388.3 km
  • Perigee — 382.0 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004641
  • Orbits per 24-hr. day — 15.60
  • Altitude decrease — 400 m (mean) in last 24 hours
  • Solar Beta Angle: -7.0 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 18895
  • Current Flight Attitude — LVLH (local vertical/local horizontal = “earth-fixed” [yaw: -10 deg, pitch: -7 deg., roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.